The number of dimes I have is 10 more than twice the number of nickels. The value of my coins are $9. How many coins do I have?
Can you write this in an equation and solve it?

Answer by nerdybill(6958)

The number of dimes I have is 10 more than twice the number of nickels. The value of my coins are $9. How many coins do I have?
Can you write this in an equation and solve it?
.
Let x = number of nickels
then
2x+10 = number of dimes
.
.05x + .10(2x+10) = 9
.05x + .20x+1 = 9
.25x+1 = 9
.25x = 8
x = 32 (nickels)
.
Dimes:
2x+10 = 2(32)+10 = 64+10 = 74